my son is in 5th grade and his home work tonight says... evaluate the exressions
when a=3 b=67 c=489 d=9570
The first problem is
1. 7c
2. 4 . c
3. 3a X b
ect.
I dont even know where to begin on how to do this!

Answer by ankor@dixie-net.com(22740)

when a=3 b=67 c=489 d=9570
The first problem is
:
1. 7c
Replace c with 489, them you have
7 * 489 = 3423
:
2. 4 . c
Replace c with 489 again
4 * 489 = 1956
:
3. 3a X b
Replace a with 3, replace b with 67
3*3*67 = 603